# Deep Learning based Visually Rich Document Content Understanding: A Survey

### v1

@@ 4.3.3. LLM-based Frameworks
LayoutLLM (Luo et al., 2024) introduces an LLM/MLLM-based approach integrated with a pretrained document understanding model to address the challenges of applying LLMs to zero-shot document understanding tasks. The input document’s visual, textual, and layout information and any question text are encoded by a pretrained LayoutLMv3 (Huang et al., 2022) encoder and projected into the same embedding space as the adopted LLM, Vicuna-7B-v1.5 (Zheng et al., 2024). The method incorporates layout-aware pretraining tasks at three levels: document-level (e.g., document summarization), region-level (e.g., layout analysis), and segment-level (e.g., MVLM). These tasks enable the model to achieve comprehensive document understanding. Additionally, a novel module called LayoutCoT is designed to help LayoutLLM focus on question-relevant regions and generate accurate answers through intermediate steps. GPT-3.5-turbo (OpenAI, 2023) is used to prepare the dataset for document summarization training and to construct LayoutCoT training data.
4.4. Non-Pretrained Frameworks
@@
CALM (Du et al., 2022) introduces a common-sense augment document understanding framework to understand the query and extrapolate answers not contained in the context of the input document image. They follow LayoutLMv2 (Xu et al., 2020b) to encode input document multimodal representations. The textual token embeddings are fed into a Document Purifier component to merge the tokens $\{t_{1},\dots,t_{n}\}$ belonging to one entity type $N$ to one Upper Layer token $\hat{c}$ by applying average pooling of $\hat{c}=AvePool(t_{1},\dots,t_{n})$. Each Upper Layer token is concatenated with the commence augmented on ConceptNet NumberBatch (Speer et al., 2017) entity word vector $c^{\prime}$ to get the final entity representation $c=concat(\hat{c},c^{\prime})$. A similar Question-Purifier is applied to use common-sense knowledge to enhance the question representation. Then, with the assistance of ConceptNet, relevant common-sense knowledge is recalled based on the common-sense representation of both documents and queries. By considering the predicted question-answer relationship, a final self-attentive graph convolutional network following (Zhu et al., 2021b) is proposed to address document reasoning tasks more effectively.
@@
LayoutGCN (Shi et al., 2023) proposes a lightweight and effective model which contains a fully connected graph where text blocks are nodes and edges connect every two blocks. The model architecture includes a TextCNN-based (Kim, 2014) encoder to encode N-gram textual embeddings, a linear trainable layout encoder to project the normalised bbox coordinates into hyperspace following other layout-aware models, and a visual encoder (CSP-Darknet (Wang et al., 2020a) for document image features). These features are integrated using a Graph Convolution Network (GCN) to capture relationships between nodes. The final node representation combines text, layout, and visual information, benefiting various VRDU tasks.
XYLayoutLM (Gu et al., 2022) introduces an Augmented XY Cut module to correct the improper reading order generated by OCR engines and a Dialted Conditional Position Encoding module to handle the variable lengths of input text and image tokens. LayoutXLM (Xu et al., 2021)  [^6] Please refer to Section 4.1.2 to see a detailed description of the LayoutXLM model. is adopted as a basic framework to process multimodal inputs. The Augmented XY Cut module enhances traditional XY Cut (Nagy and Seth, 1984) by incorporating thresholds ($\lambda_{x},\lambda_{y}$) and a shift factor ($\theta$) to adjust box positions on the x and y axis based on randomly generated values. It recursively divides token boxes using valleys in horizontal and vertical projection profiles, forming clusters in descending order. Each cluster’s reading order is determined recursively, prioritizing divisions until no significant valleys remain, ensuring a proper reading order is derived from an XY Tree structure. DCPE (Dilated Conditional Position Encoding) addresses limitations of CPE (Chu et al., 2022) in multimodal tasks by separately processing textual and visual features. It employs 1D convolutions for textual tokens to extract 1D local layouts, accommodating their inherent 1D relationships. Additionally, DCPE utilizes dilated convolutions (Yu and Koltun, 2015) to capture long-range dependencies effectively without increasing model complexity, thereby enhancing performance in multimodal document understanding tasks.
4.5. Summary of Multi-Task Frameworks
@@
Various models are proposed to enhance document representations for VRDU tasks by leveraging pretrained language models to enrich text token sequences with layout information through positional encoding (Xu et al., 2020a), attention mechanisms (Wang et al., 2022b), and layout-aware tasks (Tu et al., 2023). However, VRDs contain rich visual details like font, texture, and colour and visually complex entities such as tables, charts, and photos. Many models (Xu et al., 2020b; Huang et al., 2022; Appalaraju et al., 2021) integrate visual cues to enhance fine-grained document features, but their quadratic time and space complexity pose challenges for handling long sequences in multi-page document understanding (Ding et al., 2024a). Fine-grained models excel but struggle with capturing layout and structural details from document images. Coarse-grained frameworks (Li et al., 2021b; Gu et al., 2021) mitigate fine-grained limitations by leveraging entity-level multimodal information, yet compressing diverse entity aspects into a single dense vector risks losing information (Ding et al., 2024b). Joint-grained frameworks (Li et al., 2021c; Zhai et al., 2023; Wang et al., 2022a; Bai et al., 2022; Luo et al., 2023) integrate multi-grained information to produce comprehensive representations. Non-pretrained models leverage external knowledge (Du et al., 2022) or lightweight networks (Shi et al., 2023) to rival large-scale pretrained frameworks’ performance. Most document understanding models (Xu et al., 2020a; Huang et al., 2022; Wang et al., 2022b; Li et al., 2021b) rely on off-the-shelf OCR tools for text extraction, which can be susceptible to OCR quality issues and incorrect reading orders. OCR-free frameworks directly process document images to mitigate these limitations; However, these frameworks may exhibit sub-optimal performance compared to methods using established OCR tools with additional resource consumption.
5. Visually Rich Document Content Understanding Datasets
5.1. Key Information Extraction and Entity Linking
@@
Name Conf./J. Year Domain # Docs # Images # Keys MP. Language Metrics Format FUNSD ICDAR-w 2019 Multi-source N/A 199 4 N English F1 P.& H. SROIE ICDAR-c 2019 Scanned Receipts N/A 973 4 N English F1* P. CORD Neurips-w 2019 Scanned Receipts N/A 1,000 54 N English F1 P. Payment-Invoice ACL 2020 Invoice Form N/A 14,237+595 7 N English F1 D. Payment-Receipts ACL 2020 Scanned Receipts N/A 478 2 N English F1 P. Kleister-NDA ICDAR 2021 Private Agreements 540 3,229 4 Y English F1 D. Kleister-Charity ICDAR 2021 AFR 2,778 61,643 8 Y English F1 D.& P. EPHOIE AAAI 2021 Exam Paper N/A 1,494 10 N Chinese F1 P.& H. XFUND ACL 2022 Synthetic Forms N/A 1,393 4 N Multilingual F1 D.& P.& H. Form-NLU SIGIR 2023 Financial Form N/A 857 12 N English F1 D.& P.& H. VRDU-Regist. Form KDD 2023 Registration Form N/A 1,915 6 N English F1 D. VRDU-Ad-buy Form KDD 2023 Political Invoice Form N/A 641 9+1(5) N English F1 D.&P. DocILE ICDAR 2023 Invoice Form 6,680 106,680 55 Y English AP, CLEval D.&P.
Table 1. Summary of visually rich document key information extraction datasets.
5.1.1. Scanned Receipt Datasets
@@
SROIE (Huang et al., 2019) is a widely used dataset for text localization, OCR, and key information extraction from scanned receipts, introduced in the ICDAR 2019 Challenge on ”Scanned Receipts OCR and Key Information Extraction” The Key Information Extraction (KIE) task focuses on four key types: Address, Date, Company, and Total, with corresponding values provided in the annotation file for each receipt. The F1 score for this task is calculated based on Mean Average Precision (MAP) and recall. Note that entity-level annotations are not provided in the official dataset, requiring the use of external tools to obtain the information for coarse-grained or joint-grained models.
@@
Payment-Receipts (Majumder et al., 2020b) is a subset of SROIE, created by sampling up to 5 documents from each template in the original SROIE dataset. The template of each receipt is decided by the Company annotation. The target schema focuses on extracting only Date and Total. This subset is used to evaluate the model’s ability to handle unseen templates.
@@
CORD (Park et al., 2019) is a widely used dataset for post-OCR receipt understanding, featuring two-level labels annotated by crowdsourcing workers. It includes eight superclasses, such as Store, Payment, Menu, Subtotal, and Total, each with several subclasses. For example, Store contains subclasses like Name, Address, and Telephone. CORD provides both textline-level and word-level annotations for both fine-grained and coarse-grained frameworks, with some sensitive information blurred. All models are evaluated on the released first 1,000 samples.
5.1.2. Form-style Datasets
@@
FUNSD (Jaume et al., 2019) is derived from the RVL-CDIP dataset (Harley et al., 2015) by manually selecting 199 readable and diverse template form images. The dataset is annotated using the GuiZero library to provide both entity and word-level annotations, including manual text recognition. Semantic links indicate relationships between entities, such as Question-Answer or Header-Question pairs. Consequently, FUNSD supports key information extraction, OCR, and entity linking tasks.
@@
XFUND (Xu et al., 2021) is the first multilingual dataset following the FUNSD format. It collects form templates in seven languages (Chinese, Japanese, Spanish, French, Italian, German, and Portuguese) from the internet. Human annotators fill these templates with synthetic information by typing or handwriting, ensuring each template is used only once. The filled forms are then scanned into document images, processed with OCR, and annotated with key-value pairs. Each language has 199 annotated forms, supporting multilingual key information extraction and entity linking tasks.
@@
Payment-Invoice (Majumder et al., 2020b) contains two corpora of invoices from different sources. The first corpus, Invoice 1, includes 14,273 invoices from various vendors with different template styles, used for training and validation. The second corpus, Invoice 2, comprises 595 documents with distinct templates not found in Invoice 1, serving as the test set. Human annotators extract six required keys from each single-page invoice, such as Invoice Date, Total Amount, and Tax Amount. This dataset is suitable for evaluating generative-style models and fine-grained sequence labeling models. For coarse-grained models, additional text line or entity-level information can be extracted using off-the-shelf tools.
@@
VRDU-Registration Form (Wang et al., 2023b) is a dataset of registration forms about foreign agents registering with the US government collected from the Federal Communications Commission. Commercial OCR tools extract the text content of the forms. Annotators draw bounding boxes around six unrepeated entities (each entity appears only once per document) per document: File Date, Foreign Principal Name, Registrant Name, Registration ID, Signer Name, and Signer Title. The dataset provides entity-level annotations, which can be easily preprocessed to acquire token-level annotations, supporting any granularity of Key Information Extraction (KIE) models.
@@
VRDU-Ad-buy Form (Wang et al., 2023b) consists of 641 invoices or receipts signed between TV stations and campaign groups for political advertisements. It follows the same annotation procedure as the VRDU-Registration Form but involves a more complex schema. This includes nine unique entities (e.g., Advertiser, Agency, Contract ID), four repeated entities (e.g., Item Description, Sub Prices), and hierarchical entities (e.g., Line Item). Repeated entities may contain different values within a single document, while hierarchical entities comprise several repeated entities as components.
@@
Form-NLU (Ding et al., 2023a) is a visual-linguistics dataset designed to support researchers in interpreting specific designer intentions amidst various types of noise from different form carriers, including digital, printed, and handwritten forms. Fine-grained key-value pairs, such as Company Name, Previous Notice Date, and Previous Shares, are manually annotated. The training and validation set comprises 535 digital-born forms, with 76 reserved for validation. Additionally, three test sets are provided, containing 146 digital, 50 printed, and 50 handwritten form images, respectively. Form-NLU can be used to evaluate form layout analysis and Key Information Extraction (KIE) models of any granularity. With proper processing, it can also be used to evaluate entity linking frameworks, thanks to the well-annotated key-value pairs.
5.1.3. Multi-page Datasets
@@
Kleister-NDA (Stanisławek et al., 2021) is a dataset collected from the Electronic Data Gathering, Analysis, and Retrieval System (EDGAR) focusing on Non-disclosure Agreements (NDAs). During preprocessing, the collected 540 HTML files are converted into digital multi-page PDF files (totalling 3,229 pages) using the Puppeteer library. Four key items, Effective Date, Party, Jurisdiction, and Term, are manually annotated by three annotators to extract the corresponding values. The NDA dataset is widely used by fine-grained level models but may require additional processing for frameworks with limited sequence length due to the multi-page inputs.
@@
Kleister-Charity (Stanisławek et al., 2021) contains 2,778 annual financial reports from the Charity Commission, which lack strict formatting rules. The Charity Commission website provides eight key pieces of information, such as Postcode, Charity Name, and Report Date. Annotators manually correct minor errors to ensure accuracy. Compared to Kleister-NDA, the Charity dataset has longer document inputs, totalling 61,643 pages, requiring models to handle long sequence outputs. Both Charity and NDA datasets provide only key-value pair annotations, making them suitable for the generation and fine-grained sequence labelling tasks but requiring additional processing to acquire entity-level annotations.
@@
DocILE (Šimsa et al., 2023) comprises three subsets: an annotated set of 6,680 real business documents, an unlabeled set of 932,000 real business documents for unsupervised pretraining, and a synthetic set of 100,000 documents generated with full task labels. Documents come from public sources like the UCSF Industry Documents Library and Public Inspection Files, with annotations for Key Information Localization and Extraction and Line Item Recognition. Synthetic documents were created using annotated templates and a rule-based synthesizer. DocILE provides entity-level annotations that can be easily post-processed to acquire token-level annotations.
5.2. Visual Question Answering

### v2

@@ 4.3.3 LLM-based Frameworks
LayoutLLM [layoutllm] introduces an LLM/MLLM-based approach integrated with a pretrained document understanding model to better fuse multimodal information. The input document’s visual, textual, and layout information and any question text are encoded by a pretrained LayoutLMv3 [layoutlmv3] encoder and projected into the same embedding space as the adopted LLM, Vicuna-7B-v1.5 [zheng2024judging]. The method incorporates layout-aware pretraining tasks at three levels: document-level (e.g., document summarization), region-level (e.g., layout analysis), and segment-level (e.g., MVLM). These tasks enable the model to achieve comprehensive document understanding. Additionally, a novel module called LayoutCoT is designed to help LayoutLLM focus on question-relevant regions and generate accurate answers through intermediate steps. GPT-3.5-turbo [chatgpt] is used to prepare the dataset for document summarization training and to construct LayoutCoT training data.
HRVDA [hrvda] aims to propose a MLLM accepting high-resolution image inputs to conduct fine-grained information extraction from VRDs. A swin-transformer [liu2021swin] is used to encode document images into image patch tokens. A pluggable content detector then identifies visual tokens that contain relevant document content information. Following this, a content filtering mechanism performs token pruning to remove irrelevant tokens. The remaining encoded visual tokens are processed through an MLP to ensure consistency with the LLM embedding space dimensions. These pruned tokens are then fused with instruction features, allowing further filtering of tokens irrelevant to the instructions. The final streamlined set of visual tokens and instructions is fed into the LLM, which generates the corresponding responses.
4.4 Non-Pretrained Frameworks
@@
CALM [calm] introduces a common-sense augment document understanding framework to understand the query and extrapolate answers not contained in the context of the input document image. They follow LayoutLMv2 [layoutlmv2] to encode input document multimodal representations. The textual token embeddings are fed into a Document Purifier component to merge the tokens $\{t_{1},\dots,t_{n}\}$ belonging to one entity type $N$ to one Upper Layer token $\hat{c}$ by applying average pooling of $\hat{c}=AvePool(t_{1},\dots,t_{n})$. Each Upper Layer token is concatenated with the commence augmented on ConceptNet NumberBatch [speer2017conceptnet] entity word vector $c^{\prime}$ to get the final entity representation $c=concat(\hat{c},c^{\prime})$. A similar Question-Purifier is applied to use common-sense knowledge to enhance the question representation. Then, with the assistance of ConceptNet, relevant common-sense knowledge is recalled based on the common-sense representation of both documents and queries. By considering the predicted question-answer relationship, a final self-attentive graph convolutional network following [zhu2021mucko] is proposed to address document reasoning tasks more effectively.
@@
LayoutGCN [layoutgcn] proposes a lightweight and effective model which contains a fully connected graph where text blocks are nodes and edges connect every two blocks. The model architecture includes a TextCNN-based [textcnn] encoder to encode N-gram textual embeddings, a linear trainable layout encoder to project the normalised bbox coordinates into hyperspace following other layout-aware models, and a visual encoder (CSP-Darknet [wang2020cspnet] for document image features). These features are integrated using a Graph Convolution Network (GCN) to capture relationships between nodes. The final node representation combines text, layout, and visual information, benefiting various VRD-CU tasks.
XYLayoutLM [xylayoutlm] builds on LayoutXLM [layoutxlm] by introducing two key modules: an Augmented XY Cut and a Dilated Conditional Position Encoding (DCPE). The Augmented XY Cut improves traditional XY Cut [nagy1984hierarchical] by incorporating adjustable thresholds ($\lambda_{x},\lambda_{y}$) and a shift factor ($\theta$) to refine token clustering and correct OCR-induced reading order errors through recursive XY Tree-based segmentation. DCPE addresses the limitations of Conditional Position Encoding (CPE) [chu2022conditional] by separately modeling textual and visual modalities. It uses 1D convolutions for textual tokens to capture local sequence structure and dilated convolutions [yu2015multi] to encode long-range dependencies without added complexity, improving multimodal representation learning in document understanding tasks.
4.5 Summary of Multi-Task Frameworks
@@
Various models are proposed to enhance document representations for VRD-CU tasks by leveraging pretrained language models to enrich text token sequences with layout information through positional encoding [layoutlm], attention mechanisms [lilt], and layout-aware tasks [layoutmask]. However, VRDs contain rich visual details like font, texture, and colour and visually complex entities such as tables, charts, and photos. Many models [layoutlmv2, layoutlmv3, docformer] integrate visual cues to enhance fine-grained document features, but their quadratic time and space complexity pose challenges for handling long sequences in multi-page document understanding [mmvqa]. Fine-grained models excel but struggle with capturing layout and structural details from document images. Coarse-grained frameworks [selfdoc, udoc] mitigate fine-grained limitations by leveraging entity-level multimodal information, yet compressing diverse entity aspects into a single dense vector risks losing information [3mvrd]. Joint-grained frameworks [structext, faststructext, mgdoc, wukong, geolayoutlm] integrate multi-grained information to produce comprehensive representations. Non-pretrained models leverage external knowledge [calm] or lightweight networks [layoutgcn] to rival large-scale pretrained frameworks’ performance. Most document understanding models [layoutlm, layoutlmv3, lilt, selfdoc] rely on off-the-shelf OCR tools for text extraction, which can be susceptible to OCR quality issues and incorrect reading orders. OCR-free frameworks directly process document images to mitigate these limitations; However, these frameworks may exhibit sub-optimal performance compared to methods using established OCR tools with additional resource consumption.
5 Visually Rich Document Content Understanding Datasets
Based on the downstream tasks differences, we summarise the Key Information Extraction with Entity Linking datasets and Visually Rich Document Question Answering dataset in Section 5.1 and 5.2.
5.1 Key Information Extraction and Entity Linking
@@
Name Venue Year Domain # Docs # Images # Keys MP. Language Metrics Format FUNSD ICDAR-w 2019 Multi-source N/A 199 4 N English F1 P./ H. SROIE ICDAR-c 2019 Scanned Receipts N/A 973 4 N English F1 P. CORD Neurips-w 2019 Scanned Receipts N/A 1,000 54 N English F1 P. Payment-Invoice ACL 2020 Invoice Form N/A 14,237+595 7 N English F1 D. Payment-Receipts ACL 2020 Scanned Receipts N/A 478 2 N English F1 P. Kleister-NDA ICDAR 2021 Private Agreements 540 3,229 4 Y English F1 D. Kleister-Charity ICDAR 2021 AFR 2,778 61,643 8 Y English F1 D./ P. EPHOIE AAAI 2021 Exam Paper N/A 1,494 10 N Chinese F1 P./ H. XFUND ACL 2022 Synthetic Forms N/A 1,393 4 N Multilingual F1 D./ P./ H. Form-NLU SIGIR 2023 Financial Form N/A 857 12 N English F1 D./ P./ H. VRDU-Regist. Form KDD 2023 Registration Form N/A 1,915 6 N English F1 D. VRDU-Ad-buy Form KDD 2023 Political Invoice Form N/A 641 9+1(5) N English F1 D./P. DocILE ICDAR 2023 Invoice Form 6,680 106,680 55 Y English AP, CLEval D./P.
Table 1: Summary of key information extraction (KIE) datasets for VRDs. MP.: Multi-page; F1: F1-score; P.: Plain-text; H.: Handwritten; D.: Digital.
5.1.1 Scanned Receipt Datasets
@@
SROIE [sroie] is a widely used dataset for text localization, OCR, and key information extraction from scanned receipts, introduced in the ICDAR 2019 Challenge on ”Scanned Receipts OCR and Key Information Extraction” The Key Information Extraction (KIE) task focuses on four key types: Address, Date, Company, and Total, with corresponding values provided in the annotation file for each receipt. The F1 score for this task is calculated based on Mean Average Precision (MAP) and recall. Note that entity-level annotations are not provided in the official dataset, requiring the use of external tools to obtain the information for coarse-grained or joint-grained models.
@@
Payment-Receipts [payment] is a subset of SROIE, created by sampling up to 5 documents from each template in the original SROIE dataset. The template of each receipt is decided by the Company annotation. The target schema focuses on extracting only Date and Total. This subset is used to evaluate the model’s ability to handle unseen templates.
@@
CORD [cord] is a widely used dataset for post-OCR receipt understanding, featuring two-level labels annotated by crowdsourcing workers. It includes eight superclasses, such as Store, Payment, Menu, Subtotal, and Total, each with several subclasses. For example, Store contains subclasses like Name, Address, and Telephone. CORD provides both textline-level and word-level annotations for both fine-grained and coarse-grained frameworks, with some sensitive information blurred. All models are evaluated on the released first 1,000 samples.
5.1.2 Form-style Datasets
@@
FUNSD [funsd] is derived from the RVL-CDIP dataset [rvlcdip] by manually selecting 199 readable and diverse template form images. The dataset is annotated using the GuiZero library to provide both entity and word-level annotations, including manual text recognition. Semantic links indicate relationships between entities, such as Question-Answer or Header-Question pairs. Consequently, FUNSD supports key information extraction, OCR, and entity linking tasks.
@@
XFUND [layoutxlm] is the first multilingual dataset following the FUNSD format. It collects form templates in seven languages (Chinese, Japanese, Spanish, French, Italian, German, and Portuguese) from the internet. Human annotators fill these templates with synthetic information by typing or handwriting, ensuring each template is used only once. The filled forms are then scanned into document images, processed with OCR, and annotated with key-value pairs. Each language has 199 annotated forms, supporting multilingual key information extraction and entity linking tasks.
@@
Payment-Invoice [payment] contains two corpora of invoices from different sources. The first corpus, Invoice 1, includes 14,273 invoices from various vendors with different template styles, used for training and validation. The second corpus, Invoice 2, comprises 595 documents with distinct templates not found in Invoice 1, serving as the test set. Human annotators extract six required keys from each single-page invoice, such as Invoice Date, Total Amount, and Tax Amount. This dataset is suitable for evaluating generative-style models and fine-grained sequence labeling models. For coarse-grained models, additional text line or entity-level information can be extracted using off-the-shelf tools.
@@
VRDU-Registration Form [vrdu] is a dataset of registration forms about foreign agents registering with the US government collected from the Federal Communications Commission. Commercial OCR tools extract the text content of the forms. Annotators draw bounding boxes around six unrepeated entities (each entity appears only once per document) per document: File Date, Foreign Principal Name, Registrant Name, Registration ID, Signer Name, and Signer Title. The dataset provides entity-level annotations, which can be easily preprocessed to acquire word-level annotations, supporting any granularity of Key Information Extraction (KIE) models.
@@
VRDU-Ad-buy Form [vrdu] consists of 641 invoices or receipts signed between TV stations and campaign groups for political advertisements. It follows the same annotation procedure as the VRDU-Registration Form but involves a more complex schema. This includes nine unique entities (e.g., Advertiser, Agency, Contract ID), four repeated entities (e.g., Item Description, Sub Prices), and hierarchical entities (e.g., Line Item). Repeated entities may contain different values within a single document, while hierarchical entities comprise several repeated entities as components.
@@
Form-NLU [formnlu] is a visual-linguistics dataset designed to support researchers in interpreting specific designer intentions amidst various types of noise from different form carriers, including digital, printed, and handwritten forms. Fine-grained key-value pairs, such as Company Name, Previous Notice Date, and Previous Shares, are manually annotated. The training and validation set comprises 535 digital-born forms, with 76 reserved for validation. Additionally, three test sets are provided, containing 146 digital, 50 printed, and 50 handwritten form images, respectively. Form-NLU can be used to evaluate form layout analysis and Key Information Extraction (KIE) models of any granularity. With proper processing, it can also be used to evaluate entity linking frameworks, thanks to the well-annotated key-value pairs.
5.1.3 Multi-page Datasets
@@
Kleister-NDA [nda] is a dataset collected from the Electronic Data Gathering, Analysis, and Retrieval System (EDGAR) focusing on Non-disclosure Agreements (NDAs). During preprocessing, the collected 540 HTML files are converted into digital multi-page PDF files (totalling 3,229 pages) using the Puppeteer library. Four key items, Effective Date, Party, Jurisdiction, and Term, are manually annotated by three annotators to extract the corresponding values. The NDA dataset is widely used by fine-grained level models but may require additional processing for frameworks with limited sequence length due to the multi-page inputs.
@@
Kleister-Charity [nda] contains 2,778 annual financial reports from the Charity Commission, which lack strict formatting rules. The Charity Commission website provides eight key pieces of information, such as Postcode, Charity Name, and Report Date. Annotators manually correct minor errors to ensure accuracy. Compared to Kleister-NDA, the Charity dataset has longer document inputs, totalling 61,643 pages, requiring models to handle long sequence outputs. Both Charity and NDA datasets provide only key-value pair annotations, making them suitable for the generation and fine-grained sequence labelling tasks but requiring additional processing to acquire entity-level annotations.
@@
DocILE [docile] comprises three subsets: an annotated set of 6,680 real business documents, an unlabeled set of 932,000 real business documents for unsupervised pretraining, and a synthetic set of 100,000 documents generated with full task labels. Documents come from public sources like the UCSF Industry Documents Library and Public Inspection Files, with annotations for Key Information Localization and Extraction and Line Item Recognition. Synthetic documents were created using annotated templates and a rule-based synthesizer. DocILE provides entity-level annotations that can be easily post-processed to acquire word-level annotations.
5.1.4 Summary of KIE Datasets
KIE datasets for VRDs have progressively expanded to encompass a broader range of domains, formats, and complexities, as shown in Table 1. Early benchmarks such as FUNSD [funsd] and SROIE [sroie] primarily consist of single-page, printed documents with a limited set of predefined keys, featuring simplified annotation schemas that facilitate model prototyping and benchmarking. In contrast, recent datasets like Kleister-NDA, Kleister-Charity [nda], and DocILE [docile] incorporate multi-page documents reflective of real-world enterprise scenarios (e.g., contracts, invoices). These introduce challenges such as long context understanding, cross-page context modeling, hierarchical key structures. Further complexity arises from the inclusion of handwritten text (e.g., FUNSD, EPHOIE) and multilingual content (e.g., XFUND), necessitating models capable of handling OCR noise, variable layouts, and linguistic diversity. These trends reflect a shift toward more realistic KIE challenges, with future directions emphasizing multimodal reasoning, cross-lingual adaptability, and robust processing of complex, noisy document structures.
5.2 Visually Rich Document Question Answering